OpenStudy (anonymous):

. Solve 4 – 2q = -5(q + 1). Please show all of your work. 4+2q+5q-4=-5q+5q+ 5-4 7q=1 7q/7=1/7 q= 1/7 What did I do wrong?

OpenStudy (lalaly):

4-2q=-5q-5 4+5=2q-5q 9=-3q q=-3

OpenStudy (radar):

You erred on the first line (when you added the 5q on the left (5q-2q) does not equal 7q.